# Entropic (Gromov) Wasserstein Flow Matching with GENOT

### v3

@@ 3 Generative Entropic Neural OT
In this section, we introduce GENOT, a method to learn EOT couplings by learning their conditional distributions. In § 3.1, we focus on balanced OT, and show that GENOT approximates linear or quadratic EOT couplings, solutions to (LEOT) or (QEOT) respectively. Second, in § 3.2, we extend GENOT to the unbalanced setting by loosening the conservation of mass constraint. We define U-GENOT, which approximates solutions to (ULEOT) and (UQEOT). Finally, in § 3.3, we highlight that GENOT also adresses a fused problem, combining (LEOT) and (QEOT).
@@ 3.1 Learning Entropic Couplings with GENOT
Let $\mu\in\mathcal{M}_{1}^{+}(\mathcal{X})$, $\nu\in\mathcal{M}_{1}^{+}(\mathcal{Y})$ and $\pi^{\star}_{\varepsilon}$ be an EOT coupling between $\mu$ and $\nu$, which can be a solution of problem (LEOT) or (QEOT). By the measure disintegration theorem:
@@
Knowing $\mu$, we can hence fully describe $\pi^{\star}_{\varepsilon}$ via the conditional distributions $(\pi^{\star}_{\varepsilon}(\cdot|\mathbf{x}))_{\mathbf{x}\in\mathcal{X}}$. The latter are also of great practical interest, as they provide a way to transport a point $\mathbf{x}$ sampled from $\mu$ to the target domain $\mathcal{Y}$; either stochastically by sampling $\mathbf{y}_{1},...,\mathbf{y}_{n}$ from $\pi^{\star}_{\varepsilon}(\cdot|\mathbf{x})$, or deterministically by averaging over conditional samples: $T_{\varepsilon}(\mathbf{x}):=\mathbb{E}_{Y\sim\pi^{\star}_{\varepsilon}(\cdot|\mathbf{x})}[Y]$. Moreover, we can assess the uncertainty of these predictions using any statistic of $\pi^{\star}_{\varepsilon}(\cdot|\mathbf{x})$. This aspect is crucial for many real-world problems, especially in single-cell biology, as shown in section 5.1.
Figure 1: Sketch of the GENOT model. For each ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\mathbf{x}}$ in the support of the source ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\bm{\mu}}$, we learn a flow $\phi_{1}(\cdot|{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\mathbf{x}})$ from the noise ${\color[rgb]{0.133,0.545,0.133}\definecolor[named]{pgfstrokecolor}{rgb}{0.133,0.545,0.133}\pgfsys@color@rgb@stroke{0.133}{0.545}{0.133}\pgfsys@color@rgb@fill{0.133}{0.545}{0.133}\bm{\rho}}$ to the conditional $\pi^{\star}_{\varepsilon}(\cdot|{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\mathbf{x}})$, whose support lies in that of the target $\bm{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}\nu}}$. We highlight this procedure for a specific point ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\mathbf{x}_{1}}$.
@@
Learning the Conditional Distributions. Let $\rho=\mathcal{N}(0,I_{q})$ the standard Gaussian on the target space $\mathbb{R}^{q}\supset\mathcal{Y}$. From the noise outsourcing lemma (Kallenberg, 2002), there exists a collection of conditional generators $\{T^{\star}(\cdot|\mathbf{x})\}_{\mathbf{x}\in\mathcal{X}}$ s.t. $T^{\star}(\cdot|\mathbf{x}):\mathbb{R}^{q}\rightarrow\mathbb{R}^{q}$ and  for each $\mathbf{x}$ in the support of $\mu$, $\pi^{\star}_{\varepsilon}(\cdot|\mathbf{x})=T^{\star}(\cdot|\mathbf{x})\sharp\rho$. This means that if $Z\sim\rho$, then $Y=T^{\star}(Z|\mathbf{x})\sim\pi^{\star}_{\varepsilon}(\cdot|\mathbf{x})$. Here, we seek to learn a neural collection $\{T_{\theta}(\cdot|\mathbf{x})\}_{\mathbf{x}\in\mathcal{X}}$ of such conditional generators, fitting the constraint $T_{\theta}(\cdot|\mathbf{x})\sharp\rho=\pi^{\star}_{\varepsilon}(\cdot|\mathbf{x})$, for any $\mathbf{x}$ in the support of $\mu$. We employ the CFM framework (see § 2) and parameterize each $T_{\theta}(\cdot|\mathbf{x})$ implicitly, as the flow induced by a neural vector field $v_{t,\theta}(\cdot|\mathbf{x}):\mathbb{R}^{q}\rightarrow\mathbb{R}^{q}$. Namely, $T_{\theta}(\cdot|\mathbf{x})=\phi_{1}(\cdot|\mathbf{x}):\mathbb{R}^{q}\to\mathbb{R}^{q}$ where $\phi_{t}(\cdot|\mathbf{x})$ solves
@@
We stress that while $\mathbf{x}\in\mathcal{X}\subset\mathbb{R}^{d}$, the flow $\phi_{1}(\cdot|\mathbf{x})$ from $\rho$ to $\pi^{\star}_{\varepsilon}(\cdot|\mathbf{x})$ is defined on the target space $\mathbb{R}^{q}\supset\mathcal{Y}$. Hence, we can map samples within the same space when $p=q$, but also across incomparable spaces when $p\neq q$. In particular, this allows us to consider the quadratic OT problem (QEOT). Thus, for each $\mathbf{x}$, we optimize $v_{t,\theta}(\cdot|\mathbf{x})$ by minimizing the CFM loss (2) to map $\rho$ to $\pi^{\star}_{\varepsilon}(\cdot|\mathbf{x})$, i.e.
$\begin{split}\mathbb{E}_{t,Z\sim\rho,Y\sim\pi^{\star}_{\varepsilon}(\cdot|\mathbf{x})}[\|v_{t,\theta}\left([Z,Y]_{t}|\mathbf{x}\right)-(Y-Z)\|_{2}^{2}]\,.\end{split}$ (4)
where $[Z,Y]_{t}=(1-t)Z+tY$ interpolates between noise and conditional vectors. Averaging for all $\mathbf{x}$ in the support of $\mu$ and using Fubini’s Theorem, we derive $\mathcal{L}_{\mathrm{GENOT}}(\theta)$:
$\begin{split}\mathbb{E}_{t,Z\sim\rho,(X,Y)\sim\pi^{\star}_{\varepsilon}}[\|v_{t,\theta}\left([Y,Z]_{t}|X\right)-(Y-Z)\|_{2}^{2}]\,.\end{split}$ (5)
@@
GENOT Preserves the Coupling Information, in the sense that, in the infinite sample regime, one could provably recover the original entropic coupling, whether in the linear or quadratic OT settings, and for any cost function.
Proposition 3.1 (Well-posedness of GENOT.).
Suppose that $\mathcal{L}_{\mathrm{GENOT}}(\theta)=0$. Then, for $X\sim\mu$, $Z\sim\rho$ and $Y=\phi_{1}(Z|X)$ denoting the solution of ODE (3), one has $Y\sim\pi^{\star}_{\varepsilon}(\cdot|X)$, and hence $(X,Y)\sim\pi^{\star}_{\varepsilon}$.
This result is noteworthy since standard flow matching only preserves straight couplings (Liu et al., 2022), while bridge matching (its stochastic counterpart) only preserves the linear EOT coupling for $c(\mathbf{x},\mathbf{y})=\|\mathbf{x}-\mathbf{y}\|_{2}^{2}$, as shown recently by Bortoli et al. (2023). In contrast, GENOT is a conditional CFM model, learning the conditional distributions of the coupling independently of each other: For each $\mathbf{x}$, we leverage CFM to learn a conditional flow $\phi_{1}(\cdot|\mathbf{x})$ that maps $\rho$ to $\pi_{\varepsilon}^{\star}(\cdot|\mathbf{x})$. In this sense, GENOT can be conceptualized as an augmented flow matching procedure. We visualize that GENOT preserves the coupling information in Fig. 2.
GENOT Handles Any Cost. As GENOT preserves the coupling information, we can use it to approximate linear or quadratic EOT couplings. In both cases, we make no assumptions on the cost functions: We only need to evaluate them on samples to estimate $\pi^{\star}_{\varepsilon}$ using a discrete solver, see GENOT Alg. 1, line 5. In particular, we can use implicitly defined costs, whose evaluation requires a non-differentiable sub-routine. For instance, we can use the geodesic distance on the data manifold, which can be approximated from the shortest path distance on the $k$-nn graph induced by the Euclidean distance (Crane et al., 2013; Solomon et al., 2015). Using such a data-driven cost function in discrete OT pipelines has demonstrated its effectiveness in a range of single-cell genomics tasks (Demetci et al., 2022; Huguet et al., 2022; Klein et al., 2023), as comparing cells using $\ell_{p}$ distances often fails to capture discrepancies in gene expression space (Moon et al., 2018). With GENOT, we extend the use of this cost to the neural setting, for both linear (see § 5.1) and quadratic OT (see § 5.2).
GENOT Approximates Conditional Densities. For each $\mathbf{x}$, we build a CNF $p_{t}(\cdot|\mathbf{x})=\phi_{t}(\cdot|\mathbf{x})\sharp\rho$ between $p_{0}=\rho$ and $p_{1}=\pi^{\star}_{\varepsilon}(\cdot|\mathbf{x})$. From the instantaneous change of variables formula (Chen et al., 2018), we can then approximate the conditional density $\pi^{\star}_{\varepsilon}(\mathbf{y}|\mathbf{x})$ at an arbitrary point $\mathbf{y}\in\mathcal{Y}$. We use it, for instance, in Figures 3, 6, LABEL:fig:gw_toy_data_densities and LABEL:fig:FGW_modality_translation_cond_density.
Figure 2: Fitting the EOT coupling between two Gaussian mixtures for the Coulomb cost (Benamou et al., 2015) $c(\mathbf{x},\mathbf{y})=1/\|\mathbf{x}-\mathbf{y}\|_{2}$, and $\varepsilon=0.01$, using GENOT and OT-CFM (Tong et al., 2023b). We connect paired samples with a line. The EOT coupling pairs source samples and target samples diagonally (left). GENOT (middle) generates samples correctly, while OT-CFM (right) fails to preserve the information from mini-batch couplings. The data-setting is inspired by Bortoli et al. (2023, Fig. 1)
@@ 3.2 U-GENOT: Extension to the Unbalanced Setting
Re-Balancing the UOT Problems. In its standard form, GENOT respects marginal constraints, so it cannot directly handle unbalanced formulations (ULEOT) or (UQEOT). We show that unbalanced EOT problems can be re-balanced. Eyring et al. (2023); Lübeck et al. (2022); Yang & Uhler (2019) introduced previously these ideas in the Monge map estimation setting, namely, in a static and deterministic setup. Our method stems from the fact that, for both linear and quadratic OT, the unbalanced EOT coupling $\pi^{\star}_{\varepsilon,\tau}$ between $\mu\in\mathcal{M}^{+}(\mathcal{X})$ and $\nu\in\mathcal{M}^{+}(\mathcal{Y})$ solves a balanced EOT problem between its marginals, which are re-weighted versions $\tilde{\mu}$ and $\tilde{\nu}$ of $\mu$ and $\nu$, that have the same mass.
@@
Let $\pi^{\star}_{\varepsilon,\tau}$ be an unbalanced EOT coupling, solution of (ULEOT) or (UQEOT) between $\mu\in\mathcal{M}^{+}(\mathcal{X})$ and $\nu\in\mathcal{M}^{+}(\mathcal{Y})$. We note $\tilde{\mu}=p_{1}\sharp\pi^{\star}_{\varepsilon,\tau}$ and $\tilde{\nu}=p_{2}\sharp\pi^{\star}_{\varepsilon,\tau}$ its marginals. Then, in both cases, $\tilde{\mu}$ (resp. $\tilde{\nu}$) has a density w.r.t $\mu$ (resp. $\nu$) i.e. it exists $\eta,\xi:\mathbb{R}^{d}\rightarrow\mathbb{R}^{+}$ s.t. $\tilde{\mu}=\eta\cdot\mu$ and $\tilde{\nu}=\xi\cdot\nu$. Moreover, $\tilde{\mu}$ and $\tilde{\nu}$ have the same mass and
@@
    (Quadratic) Provided that $c_{\mathcal{X}}$ and $c_{\mathcal{Y}}$ (or $-c_{\mathcal{X}}$ and $-c_{\mathcal{Y}}$) are conditionally positive kernels (see Def. A.1), $\pi^{\star}_{\varepsilon,\tau}$ solves the balanced problem (QEOT) between $\tilde{\mu}$ and $\tilde{\nu}$ with $\varepsilon^{\prime}=m(\pi^{\star}_{\varepsilon,\tau})\,\varepsilon$, where $m(\pi^{\star}_{\varepsilon,\tau})=\pi^{\star}_{\varepsilon,\tau}(\mathcal{X}\times\mathcal{Y})$.
Learning the Coupling and the Re-Weightings Simultaneously. Thanks to Prop. 3.2, we aim to [i] learn a balanced EOT coupling between $\tilde{\mu}$ and $\tilde{\nu}$ along with [ii] the re-weighting functions $\eta,\xi$. Learning them is desirable since they model the creation and destruction of mass. We do both simultaneously by adapting the GENOT procedure. More formally, we seek to optimize $\mathcal{L}_{\mathrm{U\textrm{-}GENOT}}(\theta)$:
$\begin{split}&\mathbb{E}_{t,Z\sim\rho,(X,Y)\sim\pi^{\star}_{\varepsilon,\tau}}[\|v_{t,\theta}\left([Z,Y]_{t}|X\right)-(Y-Z)\|_{2}^{2}]\,\,\textrm{[i]}\\ &+\mathbb{E}_{X\sim\mu}[(\eta-\eta_{\theta})(X)^{2}]+\mathbb{E}_{Y\sim\nu}[(\xi-\xi_{\theta})(Y)^{2}]\,\,\,\,\,\,\,\textrm{[ii]}\end{split}$
As with GENOT, we simply need to estimate a discrete unbalanced EOT coupling $\pi_{\varepsilon,\tau}^{n}$ from samples to estimate that loss. We build upon theoretical insights on the linear OT case and extend them to the quadratic OT case in practice.
Proposition 3.3 (Poitwise estimation of re-weighting functions.).
@@
Using Prop. 3.2, $\hat{\pi}_{\varepsilon,\tau}^{n}$ is a balanced EOT coupling between its marginals, which are empirical approximations of $\tilde{\mu}$ and $\tilde{\nu}$. Hence, we estimate the term [i] of the loss as we do in the balanced case by sampling from the discrete coupling. Furthermore, Prop.3.3 highlights that the estimation of $\hat{\pi}_{\varepsilon,\tau}^{n}$ also provides a consistent estimate of the re-weighting function evaluations at each $X_{i}$ and $Y_{i}$. This enables estimating [ii]. Therefore, switching from GENOT to U-GENOT simply involves using an unbalanced solver instead of a balanced one, and regressing the re-weightings on the marginal weights of the estimated discrete coupling. We detail our procedure in Alg. 1, showing the additional steps w.r.t. GENOT in teal.

### v4

@@ 3 Generative Entropic Neural OT
We introduce GENOT, a method to learn EOT couplings (thus satisyfing N1) with any cost (N2) by learning their conditional distributions. In § 3.1, we focus on balanced OT, and show that GENOT approximates linear or quadratic EOT couplings (N4), solutions to (LEOT) or (QEOT) respectively. Then, in § 3.2, we extend GENOT to the unbalanced setting by loosening the conservation of mass constraint (N3). We define U-GENOT, which approximates solutions to (ULEOT) and (UQEOT). In § 3.3, we highlight that GENOT also adresses a fused problem, combining (LEOT) and (QEOT) (N4). Finally, we demonstrate in § 3.4 the GENOT algorithm, that allows to flexibly adapt to different OT formulations, which is key for easy usability for single-cell genomics analysts.
@@ 3.1 Learning Entropic Couplings with GENOT
Let $\mu\in\mathcal{M}_{1}^{+}(\mathcal{X})$, $\nu\in\mathcal{M}_{1}^{+}(\mathcal{Y})$ and $\pi^{\star}_{\varepsilon}$ be an EOT coupling between $\mu$ and $\nu$, which can be a solution of problem (LEOT) or (QEOT). By the measure disintegration theorem, we get
@@
Knowing $\mu$, we can hence fully describe $\pi^{\star}_{\varepsilon}$ via the conditional distributions $(\pi^{\star}_{\varepsilon}(\cdot|\mathbf{x}))_{\mathbf{x}\in\mathcal{X}}$. The latter are of great practical interest, as they provide a way to transport a point $\mathbf{x}$ sampled from $\mu$ to the target domain $\mathcal{Y}$; either stochastically by sampling $\mathbf{y}_{1},...,\mathbf{y}_{n}$ from $\pi^{\star}_{\varepsilon}(\cdot|\mathbf{x})$, or deterministically by averaging over conditional samples: $T_{\varepsilon}(\mathbf{x}):=\mathbb{E}_{Y\sim\pi^{\star}_{\varepsilon}(\cdot|\mathbf{x})}[Y]$. Moreover, we can assess the uncertainty of these predictions using any statistic of $\pi^{\star}_{\varepsilon}(\cdot|\mathbf{x})$, which is crucial in single-cell genomics.
Figure 1: Left: What do we do? One task we consider is generating RNA cell profiles from ATAC measurements and an additional cell feature. This is explained in § 5.2, and demonstrated in Fig. 4. As the cells live on manifolds in two (partially) incomparable spaces , we rely on the Fused Gromov-Wasserstein (FGW) formulation, as described in § 3.3. Here, the incomparable structural information is contained in the ATAC and the RNA measurements, while the comparable information are the cell features. Right: How do we do it? For each $({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{x}},{\color[rgb]{0.921875,0.5390625,0.17578125}\definecolor[named]{pgfstrokecolor}{rgb}{0.921875,0.5390625,0.17578125}\mathbf{u}})$ in the support of the source ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{\mu}}$, we learn a flow $\phi_{1}(\cdot|{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{x}},{\color[rgb]{0.921875,0.5390625,0.17578125}\definecolor[named]{pgfstrokecolor}{rgb}{0.921875,0.5390625,0.17578125}\mathbf{u}})$ from the noise ${\color[rgb]{0.133,0.545,0.133}\definecolor[named]{pgfstrokecolor}{rgb}{0.133,0.545,0.133}\bm{\rho}}$ to the conditional $\pi^{\star}_{\varepsilon}(\cdot|{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{x}},{\color[rgb]{0.921875,0.5390625,0.17578125}\definecolor[named]{pgfstrokecolor}{rgb}{0.921875,0.5390625,0.17578125}\mathbf{u}})$, whose support lies in that of the target $\bm{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\nu}}$. The flow is multi-modal: It allows sampling structural informations ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{y}}$, as well as features ${\color[rgb]{0.921875,0.5390625,0.17578125}\definecolor[named]{pgfstrokecolor}{rgb}{0.921875,0.5390625,0.17578125}\mathbf{v}}$ simultaneously. We highlight this procedure for a specific pair ($\mathbf{x}_{1}$, $\mathbf{u}_{1}$), with $\boldsymbol{p}=2$ and $\boldsymbol{q}=3$.
@@
Learning the Conditional Distributions. Let $\rho=\mathcal{N}(0,I_{q})$ the standard Gaussian on the target space $\mathbb{R}^{q}\supset\mathcal{Y}$. From the noise outsourcing lemma [37], there exists a collection of conditional generators $\{T^{\star}(\cdot|\mathbf{x})\}_{\mathbf{x}\in\mathcal{X}}$ s.t. $T^{\star}(\cdot|\mathbf{x}):\mathbb{R}^{q}\rightarrow\mathbb{R}^{q}$ and for each $\mathbf{x}$ in the support of $\mu$, $\pi^{\star}_{\varepsilon}(\cdot|\mathbf{x})=T^{\star}(\cdot|\mathbf{x})\sharp\rho$. This means that if $Z\sim\rho$, then $Y=T^{\star}(Z|\mathbf{x})\sim\pi^{\star}_{\varepsilon}(\cdot|\mathbf{x})$. Here, we seek to learn a neural collection $\{T_{\theta}(\cdot|\mathbf{x})\}_{\mathbf{x}\in\mathcal{X}}$ of such conditional generators, fitting the constraint $T_{\theta}(\cdot|\mathbf{x})\sharp\rho=\pi^{\star}_{\varepsilon}(\cdot|\mathbf{x})$, for any $\mathbf{x}$ in the support of $\mu$. We employ the FM framework (see § 2) and parameterize each $T_{\theta}(\cdot|\mathbf{x})$ implicitly, as the flow induced by a neural vector field $v_{t,\theta}(\cdot|\mathbf{x}):\mathbb{R}^{q}\rightarrow\mathbb{R}^{q}$. Namely, $T_{\theta}(\cdot|\mathbf{x})=\phi_{1}(\cdot|\mathbf{x}):\mathbb{R}^{q}\to\mathbb{R}^{q}$ where $\phi_{t}(\cdot|\mathbf{x})$ solves
@@
We stress that while $\mathbf{x}\in\mathcal{X}\subset\mathbb{R}^{d}$, the flow $\phi_{1}(\cdot|\mathbf{x})$ from $\rho$ to $\pi^{\star}_{\varepsilon}(\cdot|\mathbf{x})$ is defined on the target space $\mathbb{R}^{q}\supset\mathcal{Y}$. Hence, we can map samples within the same space when $p=q$, but also across incomparable spaces when $p\neq q$. In particular, this allows us to consider the quadratic OT problem (QEOT). Thus, for each $\mathbf{x}$, we optimize $v_{t,\theta}(\cdot|\mathbf{x})$ by minimizing the FM loss (2) to map $\rho$ to $\pi^{\star}_{\varepsilon}(\cdot|\mathbf{x})$, i.e.
$\begin{split}\mathbb{E}_{t,Z\sim\rho,Y\sim\pi^{\star}_{\varepsilon}(\cdot|\mathbf{x})}[\|v_{t,\theta}\left([Z,Y]_{t}|\mathbf{x}\right)-(Y-Z)\|_{2}^{2}]\,,\end{split}$ (4)
where $[Z,Y]_{t}=(1-t)Z+tY$ interpolates between noise and conditional vectors. Averaging for all $\mathbf{x}$ in the support of $\mu$ and using Fubini’s Theorem, we derive the GENOT loss:
$\begin{split}\mathcal{L}_{\mathrm{GENOT}}(\theta)\vcentcolon=\mathbb{E}_{t,Z\sim\rho,(X,Y)\sim\pi^{\star}_{\varepsilon}}[\|v_{t,\theta}\left([Y,Z]_{t}|X\right)-(Y-Z)\|_{2}^{2}]\,.\end{split}$ (5)
@@
GENOT Captures the Signal from the Mini-Batch Couplings. Standard FM only preserves straight couplings [47], while bridge matching (its stochastic counterpart) only preserves the linear EOT coupling for $c(\mathbf{x},\mathbf{y})=\|\mathbf{x}-\mathbf{y}\|_{2}^{2}$ [5]. In contrast, GENOT is a conditional FM model, learning the conditional distributions of the coupling independently of each other: For each $\mathbf{x}$, we leverage FM to learn a conditional flow $\phi_{1}(\cdot|\mathbf{x})$ that maps $\rho$ to $\pi_{\varepsilon}^{\star}(\cdot|\mathbf{x})$. Therefore, as FM preserves the marginal distributions (see 2), GENOT preserves all couplings. Thus, it always captures the signal carried out by the mini-batch couplings, in both linear and quadratic settings, and for any cost. We highlight this property in Fig. LABEL:fig:coulomb_cost, comparing GENOT to other OT-based FM approaches [85, 84, 64].
GENOT Handles Any Cost. We can use GENOT to approximate linear or quadratic EOT couplings. In both cases, we make no assumptions on the cost functions: We only need to evaluate them on samples to estimate $\pi^{\star}_{\varepsilon}$ using a discrete solver, see GENOT Alg. 1, line 5. In particular, we can use implicitly defined costs, whose evaluation requires a non-differentiable sub-routine. For instance, we can use the geodesic distance on the data manifold, which can be approximated from the shortest path distance on the $k$-nn graph induced by the Euclidean distance [12, 77]. This approach has demonstrated its effectiveness in a range of single-cell genomics tasks relying on discrete OT [17, 35, 39]. We use it with GENOT for both linear (see § 5.1) and quadratic OT (see § 5.2).
GENOT Approximates Conditional Densities. For each $\mathbf{x}$, we build a CNF $p_{t}(\cdot|\mathbf{x})=\phi_{t}(\cdot|\mathbf{x})\sharp\rho$ between $p_{0}=\rho$ and $p_{1}=\pi^{\star}_{\varepsilon}(\cdot|\mathbf{x})$. From the instantaneous change of variables formula [9], we can then approximate the conditional density $\pi^{\star}_{\varepsilon}(\mathbf{y}|\mathbf{x})$ at an arbitrary point $\mathbf{y}\in\mathcal{Y}$, proving useful to evaluate the likelihood of one-to-one matches between cells, see Fig.LABEL:fig:FGW_modality_translation_cond_density.
@@ 3.2 U-GENOT: Extension to the Unbalanced Setting
Re-Balancing the UOT Problems. In its standard form, GENOT respects marginal constraints, so it cannot directly handle unbalanced formulations (ULEOT) or (UQEOT). We show that unbalanced EOT problems can be re-balanced. Eyring et al. [20], Lübeck et al. [50], Yang and Uhler [93] introduced previously these ideas in the Monge map estimation setting, namely, in a static and deterministic setup. Our method stems from the fact that, for both linear and quadratic OT, the unbalanced EOT coupling $\pi^{\star}_{\varepsilon,\tau}$ between $\mu\in\mathcal{M}^{+}(\mathcal{X})$, $\nu\in\mathcal{M}^{+}(\mathcal{Y})$ solves a balanced EOT problem between its marginals, which are re-weighted versions $\tilde{\mu}$ and $\tilde{\nu}$ of $\mu$ and $\nu$, that have the same mass.
@@
Let $\pi^{\star}_{\varepsilon,\tau}$ be an unbalanced EOT coupling, solution of (ULEOT) or (UQEOT) between $\mu\in\mathcal{M}^{+}(\mathcal{X})$ and $\nu\in\mathcal{M}^{+}(\mathcal{Y})$. We note $\tilde{\mu}=p_{1}\sharp\pi^{\star}_{\varepsilon,\tau}$ and $\tilde{\nu}=p_{2}\sharp\pi^{\star}_{\varepsilon,\tau}$ its marginals. Then, in both cases, $\tilde{\mu}$ (resp. $\tilde{\nu}$) has a density w.r.t. $\mu$ (resp. $\nu$). That is, there exist two re-weighting functions, one on each space, $\eta:\mathcal{X}\to\mathbb{R}^{+}$ and $\xi:\mathcal{Y}\to\mathbb{R}^{+}$ s.t. $\tilde{\mu}=\eta\cdot\mu$ and $\tilde{\nu}=\xi\cdot\nu$. Furthermore, $\tilde{\mu}$ and $\tilde{\nu}$ have the same total mass and
@@
    (Quadratic) Provided that $c_{\mathcal{X}}$, $c_{\mathcal{Y}}$ (or $-c_{\mathcal{X}}$, $-c_{\mathcal{Y}}$) are CPD kernels (see Def. B.1), $\pi^{\star}_{\varepsilon,\tau}$ solves the balanced problem (QEOT) between $\tilde{\mu}$ and $\tilde{\nu}$ with $\varepsilon^{\prime}=m(\pi^{\star}_{\varepsilon,\tau})\,\varepsilon$, where $m(\pi^{\star}_{\varepsilon,\tau})=\pi^{\star}_{\varepsilon,\tau}(\mathcal{X}\times\mathcal{Y})$.
Learning the Coupling and the Re-Weightings Simultaneously. Thanks to Prop. 3.1, we aim to [i] learn a balanced EOT coupling between $\tilde{\mu}$ and $\tilde{\nu}$ along with [ii] the re-weighting functions $\eta,\xi$. Learning them is desirable since they model the creation and destruction of mass. We do both simultaneously by adapting the GENOT procedure. Formally, we seek to optimize:
$\begin{split}\mathcal{L}_{\mathrm{U\textrm{-}GENOT}}(\theta)&=\,\mathbb{E}_{t,Z\sim\rho,(X,Y)\sim\pi^{\star}_{\varepsilon,\tau}}[\|v_{t,\theta}\left([Z,Y]_{t}|X\right)-(Y-Z)\|_{2}^{2}]\,\,\textrm{[i]}\\ &+\mathbb{E}_{X\sim\mu}[(\eta-\eta_{\theta})(X)^{2}]+\mathbb{E}_{Y\sim\nu}[(\xi-\xi_{\theta})(Y)^{2}],\,\,\,\,\,\,\,\,\,\,\,\quad\textrm{[ii]}\end{split}$
where $\eta_{\theta},\xi_{\theta}$ are (non-negative) neural re-weighting functions. Crucially, similar to (balanced) GENOT, we only need to estimate a discrete unbalanced EOT coupling $\pi_{\varepsilon,\tau}^{n}$ using samples $X_{1},\dots,X_{n}\sim\mu$ and $Y_{1},\dots,Y_{n}\sim\nu$ to compute the two components, [i] and [ii], of the U-GENOT loss. We build upon theoretical insights on the linear OT case and extend them to the quadratic OT case in practice.
Proposition 3.2 (Pointwise estimation of re-weighting functions.).
@@
Using Prop. 3.1, $\hat{\pi}_{\varepsilon,\tau}^{n}$ is a balanced EOT coupling between its marginals, which are empirical approximations of $\tilde{\mu}$ and $\tilde{\nu}$. Hence, we estimate the first term [i] of the loss as we do in the balanced case by sampling from the discrete coupling. Furthermore, Prop.3.2 highlights that the estimation of $\hat{\pi}_{\varepsilon,\tau}^{n}$ also provides a consistent estimate of the re-weighting function evaluations at each $X_{i}$ and $Y_{i}$. This enables estimating the second term [ii]. Therefore, switching from GENOT to U-GENOT simply involves using an unbalanced solver instead of a balanced one, and regressing the neural re-weighting functions on the marginal weights of the estimated discrete coupling. We detail our procedure in Alg. 1, showing the additional steps w.r.t. GENOT in teal.